Suppose a 2 kg object is falling through the air and accelerating downwards at 8 m/s2. Taking g = 10 m/s2 What is the magnitude and direction of the force of air resistance on it?

2N downward

2N upward

1N downward

1N upward
Answer #1

Fnet=ma

air resistance R will exert upward and gravity will exert downward

R-mg= -ma

R=m(g-a)

R=2*2 =4N

upward

b should be correct if a=9m/m^2